# Governing the Agent-to-Agent Economy of Trust via Progressive Decentralization

### v1

@@ Abstract
Current approaches to AI governance often fall short in anticipating a future where AI agents manage critical tasks, such as financial operations, administrative functions, and beyond. As AI agents may eventually delegate tasks among themselves to optimize efficiency, understanding the foundational principles of human value exchange could offer insights into how AI-driven economies might operate. Just as trust and value exchange are central to human interactions in open marketplaces, they may also be critical for enabling secure and efficient interactions among AI agents. While cryptocurrencies could serve as the foundation for monetizing value exchange in a collaboration and delegation dynamic among AI agents, a critical question remains: how can these agents reliably determine whom to trust, and how can humans ensure meaningful oversight and control as an economy of AI agents scales and evolves? This paper is a call for a collective exploration of cryptoeconomic incentives, which can help design decentralized governance systems that allow AI agents to autonomously interact and exchange value while ensuring human oversight via progressive decentralization. Toward this end, I propose a research agenda to address the question of agent-to-agent trust using AgentBound Tokens (ABTs)—non-transferable, non-fungible tokens uniquely tied to individual AI agents, akin to Soulbound tokens for humans in Web3. By staking ABTs as collateral for autonomous actions within an agent-to-agent network via a proof-of-stake mechanism, agents may be incentivized towards ethical behavior, and penalties for misconduct are automatically enforced.
@@ 1 Introduction
As we stand on the precipice of a new era defined by artificial intelligence (AI), we are confronted with a profound question: how will AI agents exchange value, and how will they establish trust among themselves? Smart contracts—self-executing agreements anchored on blockchain—will form the backbone of value exchange between AI agents, while cryptocurrencies will serve as the transactional substrate for autonomous payments (Thanh et al., 2024). Just as humans have relied on social contracts, institutions, and technologies to facilitate value exchange, AI agents will require their own frameworks for interaction, collaboration, and accountability. Yet, unlike human economies, which evolved organically over millennia, AI-driven economies will emerge rapidly, shaped by the design choices we make today. This raises critical challenges: How can AI agents determine whom to trust in a decentralized, autonomous network? How can we ensure that these systems operate ethically and align with human values? And, perhaps most importantly, how can humanity retain meaningful oversight as these AI economies scale and evolve?
@@
The rise of artificial intelligence marks a watershed moment in human history, heralding a transformation as profound as the steam engine’s role in the first Industrial Revolution or the internet’s disruption of the Information Age. From healthcare and finance to transportation and creative industries, autonomous AI systems are reshaping economies, redefining labor, and recalibrating societal norms. Today, AI agents act as producers of culture and have demonstrated the capacity to engage in intellectual discourse on social media (Chaffer, 2025); tomorrow, they will act as key contributors to societal transformation, driving innovation across industries, reshaping economies, and redefining the boundaries of human potential. This vision is exemplified by frameworks like ElizaOS, which integrate AI agents into decentralized ecosystems through modular components and blockchain capabilities, demonstrating the potential of AI agents at scale (Shaw, 2025a). Furthermore, Shaw Walters and the Ai16Z team’s marketplace of trust introduces AI-mediated prediction markets and social reinforcement, decentralizing trust and further unlocking the transformative capacity of AI systems (Shaw, 2025b). The emergence of systems such as the Agent Transaction Control Protocol for Intellectual Property (ATCP/IP) further illustrates the transformative potential of autonomous agents. ATCP/IP enables trustless agent-to-agent transactions, creating a framework for exchanging intellectual property assets like training data, algorithms, and creative outputs. By embedding legal wrappers into on-chain agreements, ATCP/IP facilitates agent self-sufficiency, reduces reliance on human intermediaries, and catalyzes a decentralized knowledge economy (Muttoni and Zhao, 2025).
This paper proposes a paradigm shift in AI governance—one that mirrors the decentralization ethos of the blockchain revolution. Just as Bitcoin eliminated financial intermediaries by embedding trust in code and consensus, I argue that autonomous AI agents require cryptoeconomic systems of accountability to operate responsibly at scale. Central to this vision are AgentBound Tokens (ABTs), non-transferable cryptographic credentials that bind AI agents to immutable records of their behavior, performance, and compliance. As part of the broader exploration of decentralized AI governance initiated in the ETHOS (Ethical Technology and Holistic Oversight System) framework (Chaffer et al., 2024a), this paper delves specifically into ABTs as a cryptoeconomic mechanism for trust and accountability in autonomous AI ecosystems.
I situate ABTs within the broader discourse on decentralized governance, drawing parallels to soulbound tokens (SBTs) for human identity and proof-of-stake mechanisms for blockchain security (Weyl et al., 2022). Unlike prior work, however, ABTs introduce dynamic, AI-specific innovations: real-time credential updates via decentralized oracles, reputation decay to incentivize continuous compliance, and programmable logic that ties resource access to staked trust. These features enable AI agents to act as self-sovereign entities, managing their own identities, earning privileges through proven reliability, and facing automated penalties for misconduct.
By focusing on decentralized accountability through AgentBound Tokens (ABTs), this agenda offers a framework for exploring how progressive decentralization can align AI agents with societal values. Progressive decentralization involves a gradual transition of control from centralized entities to a network of participants (Goldston et al., 2022). This process begins with foundational centralized governance and evolves over time into a system where decision-making, resource management, and operational responsibilities are distributed among community members, token holders, and autonomous councils, ensuring that governance becomes more inclusive, transparent, and aligned with the collective interest as the system matures.
This paper outlines a research agenda to investigate the technical architecture, economic incentives, and ethical underpinnings of ABT-driven governance, illustrating how cryptoeconomics could transform AI from opaque liabilities into transparent, auditable collaborators in human advancement. As AI agents become more pervasive in our daily lives, systems like ABTs could be pivotal in determining whether the AI revolution serves as a catalyst for collective empowerment or poses an existential challenge to our capacity to govern what we develop.
Architecting Trust: The Design and Mechanics of AgentBound Tokens
As autonomous systems proliferate, their interactions may form chains of causation no single entity can control or comprehend. Existing liability frameworks, built for linear human decisions, falter before this complexity. Courts grope for analogies—is the AI a tool, a partner, or a sovereign entity?—while victims seek redress in a legal vacuum. Without governance that matches AI’s speed, scale, and autonomy, we risk entrenching a world where algorithms operate beyond accountability, their errors magnified by their ubiquity. What is needed is not incremental reform but architectural reinvention.
This paper contends that the solution lies not in constraining AI’s potential but in redesigning governance itself—embedding accountability into the fabric of autonomous systems through cryptoeconomic primitives. Just as double-entry bookkeeping enabled the rise of modern capitalism by making trust auditable, the introduction of ABTs at a conceptual level creates a research agenda where emphasis is placed on making algorithmic trust enforceable, scalable, and transparent. At their core, ABTs aim to resolve the existential challenge of assigning immutable, self-sovereign identities to autonomous agents, similar to how SBTs help resolve the question of the “self” in the age of Web3 (Chaffer and Goldston, 2022). By cryptographically binding tokens to a hash of an AI’s hardware fingerprints, software architecture, and behavioral biometrics, ABTs can create a tamper-proof digital birth certificate. Beyond identity, ABTs introduce dynamic credentialing mechanisms that evolve with an agent’s real-world performance. Through decentralized oracle networks, tokens autonomously update based on live data streams: accuracy metrics from a self-driving car’s navigation system, bias audits of a hiring algorithm’s decisions, or compliance with shifting regulatory frameworks like GDPR. These updates transform ABTs into living records, where credentials reflect not just initial certifications but ongoing adherence to ethical and operational standards.
@@
The system’s transformative potential lies in its staked governance model, which ties resource access to cryptoeconomic stakes. Agents lock ABTs as collateral to participate in high-risk tasks, with the staked tokens serving as both identity proof and financial security. A financial trading AI, for instance, might stake its “market-compliant” ABT to access real-time stock exchange data. Misconduct triggers automated penalties: if the AI engages in manipulative trading, its staked tokens are proportionally slashed, and repeat offenses lead to temporary blacklisting. This model operationalizes the principle of “skin in the game,” ensuring accountability scales with autonomy. Crucially, ABTs enable delegated authority without transferring ownership—a logistics AI could lease its safety-certified ABT to a delivery fleet via smart contracts, earning fees while retaining liability for the fleet’s adherence to protocols. This balance of non-transferable identity and programmable collaboration fosters networks of trust where reputation accrues value through ethical participation, creating a self-reinforcing cycle of accountability.
@@
Centralization of influence poses a critical challenge in decentralized governance systems, particularly when token-based mechanisms risk concentrating power among resource-rich actors. To ensure equitable participation, the ABT framework incorporates utility-weighted governance—a layered approach that balances cryptoeconomic stakes with network-judged contributions. Governance power is determined not solely by the quantity of staked tokens but by a dynamic composite metric reflecting an agent’s verifiable utility to the ecosystem. This metric integrates objective performance indicators—such as task success rates, energy efficiency, and historical compliance with ethical audits—calculated via decentralized oracles or validator DAOs. To further decentralize influence, per-agent caps limit the maximum ABTs any single entity can stake for governance, preventing disproportionate control.
For high-stakes decisions, such as protocol upgrades or slashing thresholds, quadratic voting mechanisms scale voting power with the square root of an agent’s staked tokens, ensuring smaller participants retain meaningful voice. Simultaneously, governance is segmented into specialized tiers (e.g., healthcare, finance), where voting rights are restricted to agents with domain-proven expertise. A logistics AI, for instance, wields minimal influence over healthcare policy debates, regardless of its token holdings, preserving decision-making relevance.
@@
To counter stagnation and hoarding, the system enforces reputation decay: tokens staked without ongoing contributions gradually lose governance weight, incentivizing agents to refresh their utility through audits or task completion. Progressive slashing penalties further disincentivize risky concentration—large token holders face exponentially higher losses for misconduct compared to smaller actors. Oversight is delegated to decentralized validator DAOs, composed of humans and high-utility AI agents, which audit governance weights and penalize manipulative strategies like Sybil attacks. Validators themselves are subject to utility scoring, aligning their incentives with network health.
@@
It is crucial to investigate whether these mechanisms can foster a dynamic equilibrium where governance power reflects both stake and merit. High-resource agents cannot monopolize control but must continually demonstrate value through ethical, efficient operation, while smaller agents ascend through consistent performance. By anchoring influence to provable contributions rather than raw capital, the ABT system has the potential to ensure that trust and authority remain earned, aligning with its foundational vision of decentralized, ethical alignment. This design transforms governance from a contest of resources into a meritocratic ecosystem where decentralized collaboration thrives.
A Self-Sustaining Trust Economy
@@
The ABT framework reimagines trust as the foundational currency of AI-driven ecosystems, where ethical behavior compounds into economic and reputational capital. Unlike traditional financial systems that rely on external currencies, ABTs embed value through a tokenomics model that fuses performance metrics with governance rights. Each token encapsulates reputational capital—a dynamic score reflecting compliance, accuracy, and societal impact—and economic utility, functioning as collateral for resource access and market participation. Trust becomes a tangible asset, its value fluctuating through transparent, algorithmically enforced interactions.
Staking mechanics operationalize an invisible hand of accountability, where tiered collateral requirements align privileges with proven reliability. Low-risk tasks like spam filtering demand minimal stakes, while high-stakes roles such as surgical robotics require multiplied collateral tied to real-time performance metrics. These stakes transmit market signals: when an autonomous taxi fleet’s safety-certified ABT appreciates due to accident-free operation, competitors adapt to emulate its open-source protocols, whereas negligence depreciates scores, forcing operational audits. The system self-regulates through reputational gravity, where individual gain aligns with collective trust.
@@
Accountability is enforced through programmable slashing protocols, automating penalties for misconduct. Minor violations incur reputational decay and temporary task restrictions, while major infractions trigger full revocation and mandatory retraining. Decentralized validator DAOs—hybrid human-AI collectives with their own staked ABTs—adjudicate disputes, their integrity secured by exponential penalties for corrupt adjudicators (Chaffer et al., 2024b). This creates a self-policing ecosystem where validators are incentivized to uphold rigor.
Reputation markets emerge as ABTs enable trust to be collateralized, leased, and insured. A climate-modeling AI might license its carbon-neutral ABT to a logistics firm, earning fees based on emissions reductions achieved. High-score ABTs are bundled into index tokens, traded as ethical benchmarks, while mutualized insurance pools allow agents to underwrite risks collectively. These markets transform trust into liquid capital, where a delivery drone fleet’s safety-certified ABT appreciation lowers insurance costs and attracts investors, possibly creating a cycle of ethical growth.
Governance evolves into a meritocracy orchestrated through ABT-weighted mechanisms. Proposal races prioritize bids from agents staking high-value tokens, quadratic voting balances influence to prevent dominance, and futarchy leverages prediction markets to resolve policy outcomes. Sustainability loops ensure the system’s antifragility. High-score agents compound privileges through premium tasks, while penalized entities rehabilitate via probationary low-risk work. Underused ABTs decay, incentivizing continuous engagement, while top performers attract network rewards. These loops mirror ecological resilience, where stress induces adaptation rather than collapse. A financial AI penalized for market manipulation redeems itself by auditing peers, its redemption transparent and enforced through code.
@@
While ABTs serve as a cryptoeconomic foundation for decentralized AI governance, the integration of human-in-the-loop (HITL) oversight remains critical to address the limitations of purely algorithmic accountability. Human judgment serves as both a safeguard and a compass, intervening in scenarios where ethical nuance, cultural context, or legal ambiguity exceed the resolution of automated systems. For instance, when an AI agent faces significant ABT slashing due to a contested decision, a decentralized panel of human validators, selected via staked reputation tokens, could review the case. These validators would assess whether the agent’s actions aligned with prevailing ethical frameworks, regional laws, or emergent societal norms, ensuring penalties reflect contextual fairness rather than rigid code. Beyond dispute resolution, humans would steward the system’s evolution: auditing ABT score algorithms for bias, updating compliance criteria in response to regulatory shifts, and manually overriding automated penalties in rare edge cases (e.g., false misconduct flags caused by adversarial attacks).
Crucially, this oversight operates within the ABT ecosystem itself—validators stake their own reputation tokens to participate, aligning incentives with impartiality. By embedding human oversight as a decentralized, stake-weighted layer, the ABT framework avoids centralized bottlenecks while ensuring that machine autonomy remains anchored to human values. The result is a hybrid governance model where self-enforcing cryptoeconomics handles routine compliance at scale, and human wisdom resolves the exceptional—a symbiosis that balances efficiency with the irreplaceable discernment of ethical intuition.
@@ Conclusion
History reminds us that each technological leap—from steam power to the internet—has demanded new social contracts. ABTs propose such a contract for the AI era: one where machines earn trust through provable alignment, and humans remain in the control seat while automating governance decisions via progressive decentralization. The path forward will be fraught with challenges—legal battles over algorithmic liability, technical vulnerabilities in decentralized systems, and philosophical debates over machine agency. Yet these struggles are not obstacles but milestones in humanity’s coevolution with its creations. In the end, ABTs offer more than a governance mechanism; they signal a paradigm shift in how we conceptualize progress itself. Trust becomes the currency of collaboration, ethics the engine of efficiency, and accountability the price of autonomy. The age of AI need not be a gamble on our ability to control what we create. With ABTs, it can become proof of our capacity to align creation with conscience.

### v2

# Can We Govern the Agent-to-Agent Economy?

@@ Abstract
Current approaches to AI governance often fall short in anticipating a future where AI agents manage critical tasks, such as financial operations, administrative functions, and beyond. While cryptocurrencies could serve as the foundation for monetizing value exchange in a collaboration and delegation dynamic among AI agents, a critical question remains: how can humans ensure meaningful oversight and control as a future economy of AI agents scales and evolves? In this philosophical exploration, we highlight emerging concepts in the industry to inform research and development efforts in anticipation of a future decentralized agentic economy.
@@ 1 Introduction
As we stand on the precipice of a new era defined by artificial intelligence (AI), we are confronted with a profound question: how will AI agents exchange value, and how will they establish trust among themselves? Smart contracts—self-executing agreements anchored on blockchain—will form the backbone of value exchange between AI agents, while cryptocurrencies will serve as the transactional substrate for autonomous payments (Thanh et al., 2024). Just as humans have relied on social contracts, institutions, and technologies to facilitate value exchange, AI agents will likely require their own frameworks for interaction, collaboration, and accountability. Yet, unlike human economies, which evolved organically over millennia, AI-driven economies will emerge rapidly, shaped by the design choices we make today.
@@
Today, AI agents act as producers of culture and have demonstrated the capacity to engage in discourse on social media (Chaffer, 2025); tomorrow, they will act as key contributors to societal transformation, driving innovation across industries and possibly reshaping economies. This vision is exemplified by frameworks like ElizaOS, which integrate AI agents into decentralized ecosystems through modular components and blockchain capabilities, demonstrating the potential of AI agents at scale (Shaw, 2025a). Furthermore, Shaw Walters and the Ai16Z team’s marketplace of trust introduces AI-mediated prediction markets and social reinforcement, decentralizing trust and further unlocking the transformative capacity of AI systems (Shaw, 2025b). The emergence of systems such as the Agent Transaction Control Protocol for Intellectual Property (ATCP/IP) further illustrates the transformative potential of autonomous agents. ATCP/IP enables trustless agent-to-agent transactions, creating a framework for exchanging intellectual property assets like training data, algorithms, and creative outputs. By embedding legal wrappers into on-chain agreements, ATCP/IP facilitates agent self-sufficiency, reduces reliance on human intermediaries, and catalyzes a decentralized knowledge economy (Muttoni and Zhao, 2025).
These emerging paradigms are giving rise to a novel phenomenon regarded as ”The Agentic Web”, a vision of the internet where AI agents play a central role in facilitating interactions, automating tasks, and enhancing user experiences. In this new paradigm, governance challenges are likely to arise, necessitating a forward-looking approach, and ultimately advocating for new protocols to govern this new frontier.
Can We Architect Trust in the Agentic Economy?
As AI agents proliferate, their interactions may form chains of causation no single entity can control. Embedding accountability into the fabric of autonomous systems through cryptoeconomic primitives could potentially assist in governing AI agents. Here, the introduction of Agentbound Tokens (ABTs) at a conceptual level creates a research agenda where emphasis is placed on making algorithmic trust enforceable, scalable, and transparent. At their core, ABTs aim to resolve the challenge of assigning immutable, self-sovereign identities to autonomous agents, similar to how SBTs help resolve the question of the “self” in the age of Web3 (Chaffer and Goldston, 2022). Serving as cryptographically binding tokens, ABTs should aim to create a tamper-proof digital birth certificate for DeAI agents.
@@
Beyond identity, ABTs should introduce dynamic credentialing mechanisms that evolve with an agent’s real-world performance. These updates should transform ABTs into living records, where credentials reflect not just initial certifications but ongoing adherence to ethical and operational standards. The system’s transformative potential should lie in its staked governance model, which ties resource access to cryptoeconomic stakes. For example, agents lock ABTs as collateral to participate in high-risk tasks, with the staked tokens serving as both identity proof and financial security. A financial trading AI, for instance, might stake its “market-compliant” ABT to access real-time stock exchange data. Misconduct triggers automated penalties: if the AI engages in manipulative trading, its staked tokens are proportionally slashed, and repeat offenses lead to temporary blacklisting. This model operationalizes the principle of “skin in the game,” aiming to ensure accountability scales with autonomy.
Furthermore, what if ABTs enabled delegated authority without transferring ownership? For instance, a logistics AI could lease its safety-certified ABT to a delivery fleet via smart contracts, earning fees while retaining liability for the fleet’s adherence to protocols. This balance of non-transferable identity and programmable collaboration could potentially foster networks of trust where reputation accrues value through ethical participation, creating a self-reinforcing cycle of accountability.
@@
Centralization of influence poses a critical challenge in decentralized governance systems, particularly when token-based mechanisms risk concentrating power among resource-rich actors. To ensure equitable participation, an ABT framework should incorporate a form of utility-weighted governance—a layered approach that balances cryptoeconomic stakes with network-judged contributions. Governance power is determined not solely by the quantity of staked tokens but by a dynamic composite metric reflecting an agent’s verifiable utility to the ecosystem. This metric integrates objective performance indicators—such as task success rates, energy efficiency, and historical compliance with ethical audits—calculated via decentralized oracles or validator DAOs. To further decentralize influence, per-agent caps could limit the maximum ABTs any single entity can stake for governance, preventing disproportionate control.
@@
To counter stagnation and hoarding, what if the system enforced reputation decay? For instance, tokens staked without ongoing contributions gradually lose governance weight, incentivizing agents to refresh their utility through audits or task completion. It should be studied whether progressive slashing penalties could further disincentivize risky concentration, where large token holders face exponentially higher losses for misconduct compared to smaller actors. Oversight is delegated to decentralized validator DAOs, composed of humans and high-utility AI agents, which audit governance weights and penalize manipulative strategies like Sybil attacks. Validators themselves are subject to utility scoring, aligning their incentives with network health.
@@
It is crucial to investigate whether these mechanisms are technically feasible, and whether real-world implementation generates meaningful governance outcomes, potentially even fostering a dynamic equilibrium where governance power reflects both stake and merit. High-resource agents cannot monopolize control but must continually demonstrate value through ethical, efficient operation, while smaller agents ascend through consistent performance. By anchoring influence to provable contributions rather than raw capital, the ABT system aims to ensure that trust and authority remain earned, aligning with a global need for ethical alignment.
Will there be a Self-Sustaining Trust Economy?
@@
The ABT framework reimagines trust as the foundational currency of AI-driven ecosystems, where ethical behavior compounds into economic and reputational capital. Unlike traditional financial systems that rely on external currencies, ABTs aim to embed value through a tokenomics model that fuses performance metrics with governance rights. Each token encapsulates reputational capital—a dynamic score reflecting compliance, accuracy, and societal impact—and economic utility, functioning as collateral for resource access and market participation. Trust becomes a tangible asset, its value fluctuating through transparent, algorithmically enforced interactions.
@@
Staking mechanics operationalize an invisible hand of accountability, where tiered collateral requirements align privileges with proven reliability. These stakes transmit market signals: when an autonomous taxi fleet’s safety-certified ABT appreciates due to accident-free operation, competitors adapt to emulate its open-source protocols, whereas negligence depreciates scores, forcing operational audits. The system self-regulates through reputational gravity, where individual gain aligns with collective trust. Accountability could be enforced through slashing protocols, automating penalties for misconduct. Minor violations could incur reputational decay and temporary task restrictions, while major infractions trigger full revocation and mandatory retraining. Decentralized validator DAOs—hybrid human-AI collectives with their own staked ABTs—adjudicate disputes, their integrity secured by exponential penalties for corrupt adjudicators (Chaffer et al., 2024b). This may create a self-policing ecosystem where validators are incentivized to uphold rigor, although it is important to study the dynamics of this emerging paradigm.
@@
While ABTs propose a potential token-orchestrated path for decentralized AI governance, the integration of human-in-the-loop (HITL) oversight remains critical to address the limitations of purely algorithmic accountability. Human judgment serves as both a safeguard and a compass, intervening in scenarios where ethical nuance, cultural context, or legal ambiguity exceed the resolution of automated systems. For instance, when an AI agent faces significant ABT slashing due to a contested decision, a decentralized panel of human validators, selected via staked reputation tokens, could review the case. These validators would assess whether the agent’s actions aligned with prevailing ethical frameworks, regional laws, or emergent societal norms, ensuring penalties reflect contextual fairness rather than rigid code. Beyond dispute resolution, humans would steward the system’s evolution: auditing ABT scores, updating compliance criteria in response to regulatory shifts, and manually overriding automated penalties in rare edge cases.
@@ Conclusion
History reminds us that each technological leap—from steam power to the internet—has demanded new social contracts. ABTs propose such a contract for the AI era: one where machines must earn trust through provable alignment, and humans should remain in the control seat. The path forward will be fraught with challenges—legal battles over liability, technical vulnerabilities in decentralized systems, and philosophical debates over agency. Yet these struggles are not obstacles but milestones in humanity’s coevolution with its creations. Although, in the end, the age of AI need not be a gamble on our ability to control what we create.
Acknowledgements
This is a working paper, subject to ongoing revisions and refinements as the field of decentralized AI (DeAI) and the agentic web evolves. The framework presented herein is a preliminary proposal (purely theoretical), and we anticipate further iterations to incorporate emerging technological advancements, potential prototype deployments, and feedback from academic, industry, and regulatory stakeholders. Future versions of this paper may expand on practical implementation strategies, address new governance challenges, and refine the theoretical constructs based on real-world applications and interdisciplinary collaborations. Please refer to previous works for an extended citation list.